Once a country is wealthy, it may be easier for it to grow quickly because of the diminishing returns to capital.

Select one:

True

False

Question text

Suppose that a country’s real output has increased. If the total number of hours worked either fell or stayed the same, then productivity has increased.

Select one:

True

False

Countries with the highest growth rates today are the ones that had the highest level of real GDP per person approximately a century ago.

Select one:

True

False

Question text

In Newland, saving amounts to $7,000 and consumption amounts to $18,000 and in Greatland, saving amounts to $9,000 and consumption amounts to $25,000. The saving rate is lower in Newland than in Greatland.

Select one:

True

False

Natural resources are the renewable and nonrenewable production inputs provided by nature and physical capital is the equipment and structures available to produce goods and services.

Select one:

True

False
Answer #1

1.

False

With diminishing return of capital, the country grows slowly, rather quickly.

---

2.

True

If total real output is increasing, but hours worked do not change, then output per hour increases. So, productivity increases.

---

3.

False

Countries who were poor, some years back, is now growing with faster rates.

---

4.

False

Working note:

Saving rate for Newland = 7000/25000 = 28%

Saving rate for Greatland = 9000/34000 = 26.47%

So, saving rate is higher in Newland, rather being lower.

---

5.

True

It is the correct differentiation between natural resources and physical capital.